Handheld fi ber-optic meters with white light po-
larization interferometry are useful for measuring
temperature, pressure, and strain in electrically
noisy environments. The fi xed costs associated
with manufacturing are $850,000 per year. If a
base unit sells for $3050 and its variable cost is
$2075, ( a ) how many units must be sold each year
for breakeven and ( b ) what will the profi t be for
sales of 3000 units per year?
Answer #1

Answer

(a)

Break even occurs when Total revenue = Total Cost or we can say that profit = Total Revenue - Total Cost = 0.

Total Cost(TC) = Total Fixed Cost + Per unit Variable cost*Q

where Total Fixed Cost = 850,000 , Per unit variable cost = 2075 and Q = Quantity

=> Total Cost = Total Fixed Cost + Per unit Variable cost*Q

=> Total Cost = 850,000 + 2075Q

Total Revenue = Price *Quantity = 3050Q

Hence, Total Revenue = Total Cost => 3050Q = 850,000 + 2075Q

=> 975Q = 850,000

=> Q = 871.795 units or (872 units approx)

(b)

Profit = Total Revenue - Total Cost

=> Profit = 3050Q - (850,000 + 2075Q)

Now it is given that Q = 3000.

Profit = Total Revenue - Total Cost = 3050*3000 - (850,000 + 2075*3000)

=> Profit = $2,075,000.